### v1

@@ 5.2 Reconciling Ensemble and Region-Specific Forecasting
Regional adaptation addresses spatial heterogeneity: different regions exhibit systematically different epidemic characteristics. The Northeast’s volatile dynamics favor models that adapt quickly (GAM), while the South’s more stable patterns allow simpler approaches (ARIMA) to perform competitively. Region-specific modeling means selecting which models to include in regional ensembles (perhaps excluding Prophet if it consistently underperforms), adjusting model hyperparameters for regional characteristics (e.g., shorter smoothing windows for volatile regions), and interpreting forecast uncertainty differently across regions based on regional predictability.
@@
The temporal performance variation we observe (9) reveals an important pattern: model performance deteriorates significantly during periods of rapid change (surges) in viral activity, suggesting potential for adaptive weighting strategies. Recent work demonstrated that adaptive weighted ensembles with exponential time decay (assigning the highest weight to the most recent performance) achieved 52% error reduction compared to static weighting approaches in contexts with irregular epidemic patterns (tsang2024adaptive). Such approaches could address the performance degradation during surge periods that is evident in our analysis.
@@
Nonetheless, several limitations must be acknowledged. This is a retrospective study, meaning forecasts were tested against past data rather than generated in real time. While helpful in evaluating model accuracy, this approach does not fully reflect the challenges of real-time forecasting, such as data delays and changing conditions, which can affect operational performance. Wastewater data quality varies across jurisdictions due to differences in WWTP size, sampling protocols, and laboratory methods. Plant-level variability in viral recovery efficiency, flow normalization, and PMMoV adjustment could contribute to regional performance differences (schenk2024sars). Our focus on phenomenological and statistical models excluded mechanistic SEIR-type models incorporating transmission dynamics and intervention effects. Hybrid approaches combining mechanistic structure with ensemble phenomenological forecasting could improve performance during periods of rapidly changing transmission rates. Our study do not explicitly stratify model performance by epidemic “phase” (rising, peak/plateau, declining). Our rolling, 10-week calibration framework is designed to mimic operational use and naturally spans multiple phases, but it does not allow formal phase-specific comparisons without additional assumptions and loss of sample size. Future work could incorporate explicit phase definitions to evaluate performance across distinct stages of epidemic dynamics.
@@ 6 Conclusion
G.C. was supported by NSF Awards 2125246 and 2026797. F.A. is supported through a 2CI Fellowship from Georgia State University.
7 Supplementary
Figure S1: National-level distributions of forecast performance across models and horizons. Boxplots show the distributions of MAE, MSE, and WIS (rows; plotted on a log scale) for 1–4 week-ahead forecasts (columns) over all national forecast dates. For each model, the vertical line inside the box marks the median, and the box width spans the interquartile range. Whiskers extend to 1.5×IQR, and gray points indicate outliers.
Figure S2: Midwest region distributions of forecast performance across models and horizons. Boxplots show the distributions of MAE, MSE, and WIS (rows; plotted on a log scale) for 1–4 week-ahead forecasts (columns) over all Midwest region forecast dates. For each model, the vertical line inside the box marks the median, and the box width spans the interquartile range. Whiskers extend to 1.5×IQR, and gray points indicate outliers.
Figure S3: Northeast region distributions of forecast performance across models and horizons. Boxplots show the distributions of MAE, MSE, and WIS (rows; plotted on a log scale) for 1–4 week-ahead forecasts (columns) over all Northeast region forecast dates. For each model, the vertical line inside the box marks the median, and the box width spans the interquartile range. Whiskers extend to 1.5×IQR, and gray points indicate outliers.
Figure S4: South region distributions of forecast performance across models and horizons. Boxplots show the distributions of MAE, MSE, and WIS (rows; plotted on a log scale) for 1–4 week-ahead forecasts (columns) over all South region forecast dates. For each model, the vertical line inside the box marks the median, and the box width spans the interquartile range. Whiskers extend to 1.5×IQR, and gray points indicate outliers.
Figure S5: West region distributions of forecast performance across models and horizons. Boxplots show the distributions of MAE, MSE, and WIS (rows; plotted on a log scale) for 1–4 week-ahead forecasts (columns) over all West region forecast dates. For each model, the vertical line inside the box marks the median, and the box width spans the interquartile range. Whiskers extend to 1.5×IQR, and gray points indicate outliers.
Figure S6: Skill scores relative to the SLR baseline at the national level across models, horizons, and metrics. Heatmaps display percent skill scores ($100\times[1-\bar{m}/\bar{m}_{\text{SLR}}]$), where $\bar{m}$ and $\bar{m}_{\text{SLR}}$ denote the average MAE, MSE, or WIS across all forecast dates for a given model and for the SLR baseline, respectively. Rows correspond to MAE, MSE, and WIS, and columns to 1–4 week-ahead horizons. Positive values (blue) indicate lower error than SLR (better performance), while values near zero (white) indicate performance similar to the baseline; SLR is therefore zero by construction. The fill scale was bounded using the 5th and 95th percentiles of the skill-score distribution, with values outside this range clipped to the endpoint colors, while cell annotations report the original percent skill values.
Figure S7: Midwest region skill scores relative to the SLR baseline across models, horizons, and metrics. Heatmaps display percent skill scores ($100\times[1-\bar{m}/\bar{m}_{\text{SLR}}]$), where $\bar{m}$ and $\bar{m}_{\text{SLR}}$ denote the average MAE, MSE, or WIS across all forecast dates for a given model and for the SLR baseline, respectively. Rows correspond to MAE, MSE, and WIS, and columns to 1–4 week-ahead horizons. Positive values (blue) indicate lower error than SLR (better performance), while values near zero (white) indicate performance similar to the baseline; SLR is therefore zero by construction. The fill scale was bounded using the 5th and 95th percentiles of the skill-score distribution, with values outside this range clipped to the endpoint colors, while cell annotations report the original percent skill values.
Figure S8: Northeast region skill scores relative to the SLR baseline across models, horizons, and metrics. Heatmaps display percent skill scores ($100\times[1-\bar{m}/\bar{m}_{\text{SLR}}]$), where $\bar{m}$ and $\bar{m}_{\text{SLR}}$ denote the average MAE, MSE, or WIS across all forecast dates for a given model and for the SLR baseline, respectively. Rows correspond to MAE, MSE, and WIS, and columns to 1–4 week-ahead horizons. Positive values (blue) indicate lower error than SLR (better performance), whereas negative values (red) indicate higher error than SLR (worse performance); by definition, SLR has a skill score of zero. To reduce the influence of extreme negative outliers on the color mapping, the fill scale was bounded using the 5th and 95th percentiles of the skill-score distribution: values below the 5th percentile and above the 95th percentile were clipped to the endpoint colors, while cell annotations report the original percent skill values.
Figure S9: South region skill scores relative to the SLR baseline across models, horizons, and metrics. Heatmaps display percent skill scores ($100\times[1-\bar{m}/\bar{m}_{\text{SLR}}]$), where $\bar{m}$ and $\bar{m}_{\text{SLR}}$ denote the average MAE, MSE, or WIS across all forecast dates for a given model and for the SLR baseline, respectively. Rows correspond to MAE, MSE, and WIS, and columns to 1–4 week-ahead horizons. Positive values (blue) indicate lower error than SLR (better performance), whereas negative values (red) indicate higher error than SLR (worse performance); by definition, SLR has a skill score of zero. To reduce the influence of extreme negative outliers on the color mapping, the fill scale was bounded using the 5th and 95th percentiles of the skill-score distribution: values below the 5th percentile and above the 95th percentile were clipped to the endpoint colors, while cell annotations report the original percent skill values.
Figure S10: West region skill scores relative to the SLR baseline across models, horizons, and metrics. Heatmaps display percent skill scores ($100\times[1-\bar{m}/\bar{m}_{\text{SLR}}]$), where $\bar{m}$ and $\bar{m}_{\text{SLR}}$ denote the average MAE, MSE, or WIS across all forecast dates for a given model and for the SLR baseline, respectively. Rows correspond to MAE, MSE, and WIS, and columns to 1–4 week-ahead horizons. Positive values (blue) indicate lower error than SLR (better performance), whereas negative values (red) indicate higher error than SLR (worse performance); by definition, SLR has a skill score of zero. To reduce the influence of extreme negative outliers on the color mapping, the fill scale was bounded using the 5th and 95th percentiles of the skill-score distribution: values below the 5th percentile and above the 95th percentile were clipped to the endpoint colors, while cell annotations report the original percent skill values.

### v2

@@ 5.1 Regional Heterogeneity in Forecast Performance
An additional consideration is that forecast performance may also have been influenced by the dominant circulating SARS-CoV-2 variant during a given period and region. Variant turnover can alter transmission dynamics, fecal shedding patterns, disease severity, and the relationship between wastewater viral activity and downstream epidemiologic indicators. Wastewater studies have shown that the association between wastewater trends and reported cases or hospitalizations can differ across dominant variant waves, with both correlation strength and lead time varying over time (zhan2023correlative; hill2023wastewater). Because our forecasting framework did not explicitly incorporate sequencing-informed variant prevalence or variant-specific covariates, these effects may have contributed to the temporal and regional differences in model performance observed here. Future work could integrate genomic surveillance information from public sources of lineage prevalence to assess whether variant-aware models improve forecast performance.
Another important limitation is that we did not explicitly adjust model comparisons for differences in surveillance-network coverage across regions or time. NWSS expanded rapidly during the pandemic, and the number, distribution, and consistency of contributing sampling sites can differ across regions and periods (adams2024national). As a result, some between-region differences in forecast performance may reflect not only underlying epidemic dynamics or model structure, but also differences in the amount of information contributing to a regional WVAL signal. Future work could evaluate forecast performance jointly with surveillance coverage metrics such as the number of contributing facilities, catchment population represented, or temporal completeness of reporting.
@@ 5.2 Reconciling Ensemble and Region-Specific Forecasting
Regional adaptation addresses spatial heterogeneity: different regions exhibit systematically different epidemic characteristics. The Northeast’s volatile dynamics favor models that adapt quickly (GAM), while the South’s more stable patterns allow simpler approaches (ARIMA) to perform competitively. However, this spatial heterogeneity may also reflect differences in wastewater surveillance infrastructure, catchment composition, and temporal coverage of contributing sites, which were not explicitly modeled in the present study. Region-specific modeling means selecting which models to include in regional ensembles (perhaps excluding Prophet if it consistently underperforms), adjusting model hyperparameters for regional characteristics (e.g., shorter smoothing windows for volatile regions), and interpreting forecast uncertainty differently across regions based on regional predictability.
@@
The temporal performance variation we observe in Figure 5 reveals an important pattern: model performance deteriorates significantly during periods of rapid change (surges) in viral activity, suggesting potential for adaptive weighting strategies. Recent work demonstrated that adaptive weighted ensembles with exponential time decay (assigning the highest weight to the most recent performance) achieved 52% error reduction compared to static weighting approaches in contexts with irregular epidemic patterns (tsang2024adaptive). Such approaches could address the performance degradation during surge periods that is evident in our analysis.
@@
Forecast performance in this study was assessed by comparing each model’s 1- to 4-week-ahead predictions against subsequently observed WVAL values that were not used in model fitting. Thus, the reported errors and uncertainty metrics reflect out-of-sample predictive performance within the retrospective rolling-origin design, rather than descriptive agreement with data used to estimate the models.
@@
Nonetheless, several limitations must be acknowledged. First, this was a retrospective forecasting evaluation, meaning that model predictions were assessed against subsequently observed historical data rather than in a live operational setting. Although this design is useful for comparing predictive performance, it does not fully capture real-time constraints such as reporting delays, data revisions, or abrupt surveillance changes. Second, our analysis was conducted on the published weekly WVAL series and did not explicitly incorporate site-level metadata, surveillance-network coverage, or sequencing-informed variant prevalence. Consequently, some regional and temporal differences in model performance may reflect differences in the number and consistency of contributing sites, variant-dependent shedding and transmission dynamics, or changes in the relationship between wastewater signals and downstream epidemiologic indicators (zhan2023correlative; adams2024national). Third, wastewater measurements may vary across jurisdictions because of differences in wastewater treatment plant characteristics, catchment composition, sampling frequency, environmental conditions, laboratory workflows, and normalization practices (bertels2022factors). These factors may influence the comparability of the observed signal across locations and over time. Finally, our modeling framework focused on phenomenological and statistical approaches and did not include mechanistic SEIR-type models or hybrid models that explicitly incorporate transmission mechanisms, interventions, or variant information. Hybrid approaches combining mechanistic structure with ensemble phenomenological forecasting could improve performance during periods of rapidly changing transmission rates.
Our study also did not explicitly stratify model performance by epidemic phase (rising, peak/plateau, declining). Our rolling 10-week calibration framework was designed to mimic operational forecasting use and naturally spans multiple phases, but it does not allow formal phase-specific comparisons without additional assumptions and loss of sample size. Future work could incorporate explicit phase definitions to evaluate performance across distinct stages of epidemic dynamics.
@@ 6 Conclusion
G.C. was supported by NSF Awards 2125246 and OISE-2412914. F.A. is supported through a 2CI Fellowship from Georgia State University.